When solving the equation 9x^2−2x=3−4(x^2−2x^2+1), Mosoeu wrote 9x^2−2x=3−4x^2+8x−4 as his first step. Which property justifies this step? [1] subtraction property of equality [2] multiplication property of equality [3] associative property of multiplication [4] distributive property of multiplication over subtraction 


1
Expert's answer
2021-05-07T11:49:38-0400

Given equation- "9^2-2x=4(x^2-2x^2+1)"


Moseu wrote- "9x^2-2x=3-4x^2+8x-4"


Here It indicates the [4] Distributive property of multiplication over subtraction.